# UNGT: Ultrasound Nasogastric Tube Dataset for Medical Image Analysis

### v1

@@ Abstract
We develop a novel ultrasound nasogastric tube (UNGT) dataset to address the lack of public nasogastric tube datasets. The UNGT dataset includes 493 images gathered from 110 patients with an average image resolution of approximately 879 $\times$ 583. Four structures, encompassing the liver, stomach, tube, and pancreas are precisely annotated. Besides, we propose a semi-supervised adaptive-weighting aggregation medical segmenter to address data limitation and imbalance concurrently. The introduced adaptive weighting approach tackles the severe unbalanced challenge by regulating the loss across varying categories as training proceeds. The presented multiscale attention aggregation block bolsters the feature representation by integrating local and global contextual information. With these, the proposed AAMS can emphasize sparse or small structures and feature enhanced representation ability. We perform extensive segmentation experiments on our UNGT dataset, and the results show that AAMS outperforms existing state-of-the-art approaches to varying extents. In addition, we conduct comprehensive classification experiments across varying state-of-the-art methods and compare their performance. The dataset and code will be available upon publication at https://github.com/NUS-Tim/UNGT.
@@ 1 Introduction
The nasogastric tube (NGT) is a type of tube inserted via the nose into the stomach for feeding and it has been widely leveraged in clinical practice. The accuracy of NGT placement is vital for patient safety because a misplaced NGT would result in food content entering the lungs, resulting in a chest infection. Conventional bedside techniques for NGT placement such as pH testing and auscultation present unique limitations [1]. For instance, the lack of gastric aspirates may limit pH testing and the transmitted sounds may interfere with the auscultation. Failure of the bedside methods will require confirmation of NGT placement using a chest X-ray [2]. However, this necessitates radiation and transport of patients to the radiology department which takes time and manpower. Ultrasound (US) has been utilized as an alternative modality to confirm the NGT placement, particularly in acute care settings like the emergency department and intensive care unit [3]. Though US scanning can provide direct tube visualization fairly quickly at the bedside, its usage presents varying challenges. For instance, the accuracy of the diagnosis highly depends on the experience, the continuous assessment is time-consuming, and varied operators may issue different diagnosis results. To this end, deep learning [4, 5] has been introduced to assist in the diagnosis. Although deep learning significantly benefits the US diagnosis, the number and coverage of structures in the US datasets for model training can be largely limited. This can be attributed to the considerable complexity of capturing US images compared with natural images [6]. This complication arises from the requirement for specialized medical equipment and the involvement of medical experts during the image acquisition process. To our knowledge, existing open-source US datasets concentrating on structures such as breast [7, 8], head [9, 10], thyroid [11], and gastrointestinal [12, 13], while there is no readily US NGT dataset.
@@
Deep learning-based medical image analysis encompasses a range of tasks such as segmentation, classification, reconstruction, and captioning. Among these, segmentation and classification have emerged as the most prevalently implemented techniques [14] due to their critical role in facilitating accurate diagnosis and treatment planning. Numerous segmentation approaches adopt UNet and its variants [4, 15, 16] as a foundational architecture. The UNet holds a symmetric design, comprising an encoder for downsampling, a decoder for upsampling, and a bottleneck layer that bridges the encoder and decoder. The incorporation of skip connections facilitates the integration of contextual and spatial information. Several approaches leverage the mean teacher and its derivation as base architecture [17, 18, 19]. The mean teacher architecture consists of a teacher and a student subnet, in which the parameter of the teacher network is updated from the student network through the exponential moving average algorithm. Besides, some methods adopt the architecture with a single encoder paired with multiple decoders [20, 21, 22] and leverage varying loss function designs to adapt to diverse segmentation targets. The latter two categories are particularly prevalent in semi-supervised learning scenarios, which enables models to be trained effectively with limited data. Despite the advancements in reducing data requirements, these approaches are not specifically designed to address data imbalance. While several studies have attempted to mitigate the issue of unbalanced data [23, 24], they can overlook the concurrent challenge posed by limited data. Concerning classification, a significant portion of research leverages ResNet and its variants [25, 26, 27] as the foundation. The ResNet introduces the residual connection and has notably mitigated the difficulties associated with training deep networks thus enabling the model to capture rich and discriminative features.
@@
To bridge the dataset gap in the context of US NGT research, we introduce a novel US NGT (UNGT) dataset for medical image analysis. The UNGT dataset comprises 493 US images collected from 110 patients with an average image resolution of around 879 $\times$ 583. Structures including the liver, stomach, tube, and pancreas are precisely annotated. To concurrently address data limitation and imbalance, we introduce a semi-supervised adaptive-weighting aggregation medical segmenter (AAMS). The proposed adaptive weighting (ADW) method resolves the severe unbalanced challenge by regulating the loss across varying categories during training. The introduced multiscale attention aggregation (MAA) block enhances the feature representation by incorporating local and global contextual information. Accordingly, the AAMS can prioritize infrequent or minor structures and maintain superior representation ability. The flexibility to different unbalanced data and superior representation capability ensure the adaptability and variability of the AAMS. We perform extensive segmentation experiments leveraging AAMS and varying state-of-the-art (SOTA) approaches and the results demonstrated that the proposed AAMS shows apparent performance leadership. Moreover, we perform comprehensive classification experiments on different SOTA methods and compare their performance. To sum up, our main contributions are:
@@
    We conduct comprehensive segmentation experiments and the proposed model outperforms prevalent state-of-the-art methods. Besides, we perform extensive classification studies on varying approaches and evaluate their performance.
The rest of this paper is organized as follows. Section 2 ”Related Work” demonstrates the recent progress of the medical US dataset and medical image analysis, including segmentation and classification. Section 3 ”Methods” discusses the development of the UNGT dataset and AAMS in detail. In Section 4 ”Experiments”, the data preprocessing pipeline, experimental configuration, and evaluation metrics are presented. Detailed experimental results and analysis on the UNGT dataset can be found in Section 5 ”Results and Analysis”. We conclude our research and point out the future perspectives in Section 6 ”Conclusion”.
@@ 2.1 Medical Ultrasound Dataset
Numerous US datasets have been developed for medical image analysis on varying structures, such as breast, head, thyroid, and gastrointestinal. For instance, Yap et al. [7, 28] developed a BUS dataset collected from the UDIAT Diagnostic Centre of the Parc Taulí Corporation leveraging the Siemens ACUSON Sequoia C512 system. It comprises 163 images with an average resolution of 760 $\times$ 570, where 110 feature benign lesions and 53 feature cancerous masses. Al et al. [8] proposed a BUSI dataset collected from 600 female patients aged between 25 and 75 years using the LOGIQ E9 and LOGIQ E9 Agile US systems. The BUSI comprises 780 images with an average resolution of approximately 500 $\times$ 500, in which 437, 210, and 133 are benign, malignant, and normal. Pedraza et al. [11, 29] constructed a DDTI dataset with 637 B-mode thyroid US images curated at the IDIME US Department in Columbia. The DDTI comprises a variety of lesions such as thyroiditis, cystic nodules, adenomas, and thyroid cancers. Lu et al. [10] proposed an intrapartum transperineal JNU-IFM dataset of the Intelligent Fetal Monitoring Lab of Jinan University. It comprises 6224 US images extracted from 78 videos of 51 pregnant women and encompasses four categories depending on symphysis pubis and fetal head availability. Leclerc et al. [30] developed a CAMUS dataset with 2000 US images from 500 patients. The images were collected at the University Hospital of St Etienne and annotated in end-systole and end-diastole frames on the left ventricle endocardium, myocardium, and left atrium. Additionally, Alzubaidi et al. [9] proposed a large-scale dataset for fetal head biometry with 3832 US images. The dataset highlights the brain, cavum septum pellucidum, and lateral ventricles and the dimension of the images is 959 $\times$ 661. In 2024, Belghith et al. [31] constructed a plantar flexor muscle dataset comprising 48 images using an Aixplorer scanner in Shear Wave mode. Half of the images were captured in the prone position while the remaining were in the quiet standing position.
@@
Several medical datasets have focused on stomach-related US imaging. For example, He et al. [12] developed a GIST514-DB dataset with 251 leiomyoma images and 263 gastrointestinal stromal tumor lesion images collected in the endoscopy center of the General Hospital of Tianjin Medical University. Minoda et al. [32] collected 3980 endoscopic US images from 273 patients leveraging conventional echoendoscopes or mini-probes. In 2023, Jha et al. [13] proposed a multi-center gastrointestinal endoscopy dataset termed GastroVision. The GastroVision includes 27 anatomical landmarks and 8000 images collected from Bærum Hospital in Norway and Karolinska University Hospital in Sweden. In addition, Hirai et al. [33] collected 16110 upper gastrointestinal images from 631 cases at varying hospitals such as the Nagoya University Hospital. The subepithelial lesions include gastrointestinal stromal tumor, leiomyoma, schwannoma, neuroendocrine tumor, and ectopic pancreas. The frequency was set to 12 or 20 MHz or 5 to 7.5 MHz should the entire image cannot be captured. Despite the substantial efforts by researchers to create stomach-related US datasets, there remains a discernible void in datasets tailored for NGT investigations. To address the gap in research endeavors, we introduce a novel UNGT dataset.
@@ 2.2 Medical Image Analysis
Segmentation is one of the most significant tasks in medical image analysis. A plethora of approaches adopt UNet and its variants as the foundation. For instance, Zhou et al. [15] developed a UNet ++ approach, in which the encoder and decoder are connected through nested dense skip pathways. Subsequent work is seen in the UNet 3+ [34] with full-scale skip connections and deep supervision. Cao et al. [35] constructed a Swin-Unet model, where the tokenized image patches are fed into the transformer-based encoder-decoder to capture local and global semantic features. Chen et al. [16] developed a TransUNet, in which the ViT [36] encodes tokenized patches from convolutional. The decoder upsamples the encoded features followed by the integration with convolutional feature maps. Valanarasu [37] et al. proposed a UNeXt with an early convolutional stage and a multilayer perceptron stage, in which the tokenized multilayer perception block tokenizes and projects the convolutional features.
@@
Several methods leverage the mean teacher and its derivation as base architecture. Yu et al. [17] introduced a UA-MT approach where the teacher model generates target outputs along with the uncertainty of each prediction using Monte Carlo sampling. Wang et al. [19] constructed an MCF approach and leveraged a contrastive difference review module to alternative the potential moving average. A rectification loss grounded in the ratio of the potential mispredicted area loss with the area size was also proposed. Besides, some approaches adopt the architecture with a single encoder paired with multiple decoders and leverage varied loss functions to adapt to diverse segmentation targets. For example, Wu et al. [20] proposed an MC-Net + with a shared encoder and multiple varying decoders. The proposed method leverages the statistical discrepancy of the decoder outputs to denote the model uncertainty and introduce a mutual consistency constraint across the probability output and soft pseudo labels for varying decoders. In 2024, we [22] introduced a DEMS model that features an encoder-decoder architecture and incorporates the online automatic augmenter (OAA) and residual robustness enhancement (RRE) blocks. These approaches leverage the semi-supervised learning scheme, enabling models to be trained effectively with limited data. Despite the advancements in reducing data requirements, these approaches are not tailored to address data imbalance. Several studies have attempted to mitigate the issue of unbalanced data. For example, Pan [23] et al. proposed a cost-sensitive learning strategy to solve the imbalanced data distribution by increasing the penalization for the error classification of the minority classes. Wang [24] et al. developed a distribution-aware debiased weighting scheme and a difficulty-aware debiased weighting scheme to guide the model to solve biases. However, these approaches can overlook the concurrent challenge posed by limited data. To jointly handle data limitation and imbalance, we propose a novel AAMS model.
@@
Aside from segmentation, medical image classification serves as an important application scenario. For example, Wang et al. [5] introduced a COVID-Net with high architectural diversity and selective long-range connectivity. The intensive utilization of the projection-expansion-projection pattern ensures boosted representational capacity while maintaining computing efficiency. Lu et al. [26] proposed a revised ResNet based on the pyramid dilated convolution for Gliomas classification. Specifically, the pyramid dilated convolution was integrated into the bottom of ResNet to achieve a higher receptive field. Sharma et al. [27] constructed an enhanced ResNet by removing the final layer and adding four additional layers for brain tumor detection. Aladhadh et al. [38] leveraged the ViT to recognize skin images for skin cancer diagnosis. Images were processed with data augmentation techniques like rotation, flip, contrast, and scaling. In addition, Liang [39] et al. developed a hybrid lightweight model architecture by connecting convolution layers and MobileViT blocks [40] in tandem. In 2024, Song et al. [41] introduced a novel worse-case boosting learning algorithm to classify cervical cell images in under-representative cases. The worse-case data was sampled per gradient norm information and the enhanced loss values are then leveraged to update the classifier. Jiménez et al. [42] leveraged various generative adversarial networks to synthesize US and mammography mass images, enhancing the classification performance of ResNet. Huang et al. [43] developed an SRT approach that integrates residual blocks and triplet loss into the swin transformer [44]. This integration enhances the sensitivity to thyroid nodule features and improves the ability to distinguish subtle feature differences.

### v2

@@ Abstract
We develop a novel ultrasound nasogastric tube (UNGT) dataset to address the lack of public nasogastric tube datasets. The UNGT dataset includes 493 images gathered from 110 patients with an average image resolution of approximately 879 $\times$ 583. Four structures, encompassing the liver, stomach, tube, and pancreas, are precisely annotated. Besides, we propose a semi-supervised adaptive-weighting aggregation medical segmenter to address data limitation and imbalance concurrently. The introduced adaptive weighting approach tackles the severe unbalanced challenge by regulating the loss across varying categories as training proceeds. The presented multiscale attention aggregation block bolsters the feature representation by integrating local and global contextual information. With these, the proposed AAMS can emphasize sparse or small structures and feature enhanced representation ability. We perform extensive segmentation experiments on our UNGT dataset, and the results show that AAMS outperforms existing state-of-the-art approaches to varying extents. In addition, we conduct comprehensive classification experiments across varying state-of-the-art methods and compare their performance. The dataset and code will be available upon publication at https://github.com/NUS-Tim/UNGT.
@@ 1 Introduction
The nasogastric tube (NGT) is a type of tube inserted via the nose into the stomach for feeding, and it has been widely leveraged in clinical practice. The accuracy of NGT placement is vital for patient safety because a misplaced NGT would result in food content entering the lungs, resulting in a chest infection. Conventional bedside techniques such as pH testing and auscultation present unique limitations [13]. For instance, the lack of gastric aspirates may limit pH testing, and the transmitted sounds may interfere with the auscultation. Failure of the bedside methods will require confirmation of NGT placement using a chest X-ray [66]. However, this necessitates radiation and transport of patients to the radiology department, which takes time and manpower. Ultrasound (US) has been utilized as an alternative modality to confirm the NGT placement, particularly in acute care settings like the emergency department and intensive care unit [24]. Though US scanning can provide direct tube visualization fairly quickly at the bedside, its usage presents varying challenges. For instance, the accuracy of the diagnosis highly depends on the experience, the continuous assessment is time-consuming, and different operators may issue different diagnosis results. To this end, deep learning [44, 58] has been introduced to assist in the diagnosis. Although deep learning significantly benefits the US diagnosis, the number and coverage of structures in the US datasets for model training can be largely limited. This can be attributed to the considerable complexity of capturing US images compared with natural images [10]. This complication arises from the requirement for specialized medical equipment and the involvement of medical experts during the image acquisition process. To our knowledge, existing open-source US datasets concentrating on structures such as breast [64, 1], head [3, 32], thyroid [14], and gastrointestinal [18, 22], while there is no readily available US NGT dataset.
@@
Deep learning-based medical image analysis encompasses a range of tasks [29, 8] such as segmentation, classification, detection, reconstruction, and captioning. Among these, segmentation and classification have emerged as the most prevalently implemented techniques [72, 62, 9, 71] due to their critical role in facilitating accurate diagnosis and treatment planning. Numerous segmentation approaches adopt UNet and its variants [44, 70, 6] as a foundational architecture. The UNet holds a symmetric design, comprising an encoder for downsampling, a decoder for upsampling, and a bottleneck layer that bridges the encoder and decoder. The incorporation of skip connections facilitates the integration of contextual and spatial information. Several approaches leverage the mean teacher and its derivation as base architecture [67, 69, 59]. The mean teacher architecture consists of a teacher and a student subnet, in which the parameter of the teacher network is updated from the student network through the exponential moving average algorithm. Besides, some methods adopt the architecture with a single encoder paired with multiple decoders [63, 50, 28] and leverage varying loss function designs to adapt to diverse segmentation targets. The latter two categories are particularly prevalent in semi-supervised learning scenarios, which enables models to be trained effectively with limited data. Despite the advancements in reducing data requirements, these approaches are not specifically designed to address data imbalance. While several studies have attempted to mitigate the issue of unbalanced data [40, 57], they can overlook the concurrent challenge posed by limited data. Concerning classification, a significant portion of research leverages ResNet and its variants [17, 33, 46] as the foundation. The ResNet introduces the residual connection and has notably mitigated the difficulties associated with training deep networks, thus enabling the model to capture rich and discriminative features.
@@
To bridge the dataset gap in the context of US NGT research, we introduce a novel US NGT (UNGT) dataset for medical image analysis. The UNGT dataset comprises 493 US images collected from 110 patients with an average image resolution of around 879 $\times$ 583. Structures including the liver, stomach, tube, and pancreas are precisely annotated. To concurrently address data limitation and imbalance, we introduce a semi-supervised adaptive-weighting aggregation medical segmenter (AAMS). The proposed adaptive weighting (ADW) scheme resolves the severe unbalanced challenge by regulating the loss across varying categories during training. The introduced multiscale attention aggregation (MAA) block enhances the feature representation by incorporating local and global contextual information. Accordingly, the AAMS can prioritize infrequent or minor structures and maintain superior representation ability. The flexibility to handle different unbalanced data and superior representation capability ensure the adaptability and variability of the AAMS. We perform extensive segmentation experiments leveraging AAMS and varying state-of-the-art (SOTA) approaches, and the results demonstrate that the proposed AAMS shows apparent performance leadership. Moreover, we perform comprehensive classification experiments on different SOTA methods and compare their performance. To sum up, our main contributions are:
@@
    We conduct comprehensive segmentation experiments, and the proposed model outperforms prevalent state-of-the-art methods. Besides, we perform extensive classification studies on varying approaches and evaluate their performance.
The rest of this paper is organized as follows. Section 2 ”Related Work” demonstrates the recent progress of medical US datasets and medical image analysis, including segmentation and classification. Section 3 ”Methods” discusses the development of the UNGT dataset and AAMS model in detail. In Section 4 ”Experiments”, the data preprocessing pipeline, experimental configuration, and evaluation metrics are presented. Detailed experimental results and analysis on the UNGT dataset are in Section 5 ”Results and Analysis”. Section 6 ”Ablation and Generalization” presents the ablation study of AAMS and evaluates its generalization ability on an open-source dataset. We conclude our research and outline future perspectives in Section 7 ”Conclusion”.
@@ 2.1 Medical Ultrasound Dataset
Numerous US datasets have been developed for medical image analysis on varying structures, such as breast, head, thyroid, and gastrointestinal. For instance, Yap et al. [64, 65] developed a BUS dataset collected from the UDIAT Diagnostic Centre of the Parc Taulí Corporation using the Siemens ACUSON Sequoia C512 system. It comprises 163 images with an average resolution of 760 $\times$ 570, where 110 feature benign lesions and 53 feature cancerous masses. Al et al. [1] proposed a BUSI dataset collected from 600 female patients aged between 25 and 75 years using the LOGIQ E9 and LOGIQ E9 Agile US systems. The BUSI comprises 780 images with an average resolution of approximately 500 $\times$ 500, in which 437, 210, and 133 are benign, malignant, and normal. Pedraza et al. [14, 41] constructed a DDTI dataset with 637 B-mode thyroid US images curated at the IDIME US Department in Columbia. The DDTI comprises a variety of lesions such as thyroiditis, cystic nodules, adenomas, and thyroid cancers. Lu et al. [32] proposed an intrapartum transperineal JNU-IFM dataset of the Intelligent Fetal Monitoring Lab of Jinan University. It comprises 6224 US images extracted from 78 videos of 51 pregnant women and encompasses four categories depending on symphysis pubis and fetal head availability. Leclerc et al. [25] developed a CAMUS dataset with 2000 US images from 500 patients. The images were collected at the University Hospital of St Etienne and annotated in end-systole and end-diastole frames on the left ventricle endocardium, myocardium, and left atrium. Additionally, Alzubaidi et al. [3] proposed a large-scale dataset for fetal head biometry with 3832 US images. The dataset highlights the brain, cavum septum pellucidum, and lateral ventricles, and the dimension of the images is 959 $\times$ 661. In 2024, Belghith et al. [4] constructed a plantar flexor muscle dataset comprising 48 images using an Aixplorer scanner in Shear Wave mode. Half of the images were captured in the prone position, while the remaining were in the quiet standing position.
@@
Several medical datasets have focused on stomach-related US imaging. For example, He et al. [18] developed a GIST514-DB dataset with 251 leiomyoma images and 263 gastrointestinal stromal tumor lesion images collected in the endoscopy center of the General Hospital of Tianjin Medical University. Minoda et al. [38] collected 3980 endoscopic US images from 273 patients with conventional echoendoscopes or mini-probes. In 2023, Jha et al. [22] proposed a multi-center gastrointestinal endoscopy dataset termed GastroVision. The GastroVision includes 27 anatomical landmarks and 8000 images collected from Bærum Hospital in Norway and Karolinska University Hospital in Sweden. In addition, Hirai et al. [19] collected 16110 upper gastrointestinal images from 631 cases at varying hospitals such as the Nagoya University Hospital. The subepithelial lesions include gastrointestinal stromal tumor, leiomyoma, schwannoma, neuroendocrine tumor, and ectopic pancreas. The frequency was set to 12 or 20 MHz, or 5 to 7.5 MHz when the entire image could not be captured. Despite the substantial efforts by researchers to create stomach-related US datasets, there remains a discernible void in datasets tailored for NGT investigations. To address the gap in research endeavors, we introduce a novel UNGT dataset.
@@ 2.2 Medical Image Analysis
Segmentation is one of the most significant tasks in medical image analysis. A plethora of approaches adopt UNet and its variants as the foundation. For instance, Zhou et al. [70] developed a UNet++ approach, in which the encoder and decoder are connected through nested dense skip pathways. Subsequent work is seen in the UNet 3+ [20] with full-scale skip connections and deep supervision. Cao et al. [5] constructed a Swin-Unet model, where the tokenized image patches are fed into the transformer-based encoder-decoder to capture local and global semantic features. Chen et al. [6] developed a TransUNet that leverages ViT [12] to encode tokenized patches. The decoder upsamples the encoded features, followed by the integration with convolutional feature maps. Valanarasu [53] et al. proposed a UNeXt with an early convolutional stage and a multilayer perceptron stage, in which the tokenized multilayer perceptron block tokenizes and projects the convolutional features. Tagnamas et al. [48] developed a SCA-InceptionUNeXt that integrates a modified InceptionNeXt block and a spatial-aware channel attention module into the U-shaped architecture. Ma et al. [36] introduced a Semi-Mamba-UNet to integrate a Mamba-based UNet architecture with a conventional convolution-based UNet into the semi-supervised learning framework. Both networks generate pseudo-labels and perform mutual pixel-level supervision. Dong et al. [11] proposed an uncertainty-aware consistency learning method to improve generalization and mitigate suboptimal performance in semi-supervised segmentation.
@@
Several methods leverage the mean teacher and its derivation as the base. Yu et al. [67] introduced a UA-MT approach where the teacher model generates target outputs along with the uncertainty of each prediction using Monte Carlo sampling. Wang et al. [59] constructed an MCF approach and leveraged a contrastive difference review module to alternative the potential moving average. A rectification loss grounded in the ratio of the potential mispredicted area loss to the area size was also proposed. Wang et al. [60] introduced a S4CVNet, which comprises a feature-learning module and a guidance module. The feature-learning module employs a dual-view co-training strategy that leverages the advantages of the convolutional network and ViT while mitigating architectural discrepancies. Besides, some approaches adopt the architecture with a single encoder paired with multiple decoders and leverage varied loss functions to adapt to diverse segmentation targets. For example, Wu et al. [63] proposed an MC-Net+ with a shared encoder and multiple varying decoders. The proposed method leverages the statistical discrepancy of the decoder outputs to denote model uncertainty. Besides, it introduces a mutual consistency constraint across the probability output and soft pseudo labels across varying decoders. In 2024, we [28] introduced a DEMS model that features an encoder-decoder architecture and incorporates the online automatic augmenter (OAA) and residual robustness enhancement (RRE) blocks. These approaches leverage the semi-supervised learning scheme, enabling models to be trained effectively with limited data. Despite the advancements in reducing data requirements, these approaches are not tailored to address data imbalance. Several studies have attempted to mitigate the issue of unbalanced data. For example, Pan [40] et al. proposed a cost-sensitive learning strategy to solve the imbalanced data distribution by increasing the penalization for the error classification of the minority classes. Wang [57] et al. developed a distribution-aware debiased weighting scheme and a difficulty-aware debiased weighting scheme to guide the model to solve biases. However, these approaches can overlook the concurrent challenge posed by limited data. To jointly handle data limitation and imbalance, we propose a novel AAMS model.
@@
Aside from segmentation, medical image classification serves as an important application scenario. For example, Wang et al. [58] introduced a COVID-Net with high architectural diversity and selective long-range connectivity. The intensive utilization of the projection-expansion-projection pattern ensures boosted representational capacity while maintaining computing efficiency. Lu et al. [33] proposed a revised ResNet based on the pyramid dilated convolution for Gliomas classification. Specifically, the pyramid dilated convolution was integrated into the bottom of ResNet to achieve a higher receptive field. Sharma et al. [46] constructed an enhanced ResNet by removing the final layer and adding four additional layers for brain tumor detection. Aladhadh et al. [2] leveraged the ViT to recognize skin images for skin cancer diagnosis. Images were processed with data augmentation techniques like rotation, flip, contrast, and scaling. In addition, Liang [26] et al. developed a hybrid lightweight model architecture by connecting convolution layers and MobileViT blocks [37] in tandem. In 2024, Song et al. [47] introduced a novel worse-case boosting learning algorithm to classify cervical cell images in under-representative cases. The worse-case data was sampled per gradient norm information, and the enhanced loss values were then leveraged to update the classifier. Jiménez et al. [23] leveraged various generative adversarial networks to synthesize US and mammography mass images, enhancing the classification performance of ResNet. Huang et al. [21] developed an SRT approach that integrates residual blocks and triplet loss into the SwT [27]. This integration enhances the sensitivity to thyroid nodule features and improves the ability to distinguish subtle feature differences.